OpenStudy (anonymous):

a woman rents a car at a company that charges 30000 per day and 150 per mile travel, if the woman pays 108000 per 2 days, find the ecuation that express the amount of traveled miles by the woman. Please some help on with this?

OpenStudy (hoblos):

let d represents the number of days and m represents the number of miles and t represents the total payment so t = 30000d + 150m 150m = t - 30000d m = (t/150) - 200d for t = 108000 and d=2 m= (108000/150) - 200(2) = 720 - 400 = 320 miles

OpenStudy (anonymous):

thank my friend, just one more question this would be the ecuation for the problem t = 30000d + 150m?

OpenStudy (hoblos):

equation that express the amount of traveled miles would be m = (t/150) - 200d

OpenStudy (anonymous):

alright thanks.

OpenStudy (hoblos):

any time :)